OBSERVATION OF HIGHLY EXCITED VIBRATIONAL LEVELS OF SiH_{2} BY STIMULATED EMISSION PUMPING SPECTROSCOPY

\underline{H. ISHIKAWA}, Y. MURAMOTO, and N. MIKAMI, Department of Chemistry, Graduate School of Science, Tohoku University, Sendai 980-8578, Japan.

SiH_{2} is one of the simplest silicon containing radicals. Similarity and dissimilarity to CH_{2} radical are very interesting from both chemical and spectroscopic points of view. Renner-Teller interaction between the Ã ^{1}B_{1} and the X̃ ^{1}A_{1} states and an energy separation between the X̃ ^{1}A_{1} and the ã ^{3}B_{1} states are very important issues in spectroscopy. An analysis of highly excited vibrational levels in the X̃ state should provide some valuable information about them. Thus, we have carried out stimulated emission pumping (SEP) spectroscopy on SiH_{2}. In the present study, a pulsed discharge nozzle was used to generate jet-cooled SiH_{2} radical. Precursor of SiH_{2} radical was phenylsilane. The 0_{00} rotational levels of the Ã(060) and (070) levels were used as intermediate levels in the present SEP spectroscopy. Until now, vibrational levels with v_{2} up to 4 (E_{vib} ≤ 4000 cm^{−1}) have been observed ^{a}. In the present study, 22 vibrational levels in the energy region of 4800 – 8000 cm^{−1} were observed. Due to strong 2ν_{2}:1ν_{1} Fermi and 2ν_{1}:2ν_{3} Darling-Dennison resonances, vibrational levels having the same polyad quantum number, P = 2v_{1} + v_{2} + 2v_{3}, construct polyad structures. The vibrational levels observed belong to polyads of P = 5 − 8. Taking these resonance into account, vibrational analysis of the SiH_{2} radical was carried out. In the present analysis, vibrational constants, x_{13} and x_{23}, could be determined experimentally, for the first time. Preliminary values are -74.83 cm^{−1} and -17.58 cm^{−1}, respectively. Details of the vibrational analysis will be discussed at the presentation.

aH. ISHIKAWA and O. KAJIMOTO, J. Mol. Spectrosc. 174, 270 (1995).